# TDAD: Test-Driven Agentic Development – Reducing Code Regressions in AI Coding Agents via Graph-Based Impact Analysis

### 18 Mar 2026 v1

@@ Abstract.
AI coding agents can resolve real-world software issues, yet they frequently introduce regressions, breaking tests that previously passed. Current benchmarks focus almost exclusively on resolution rate, leaving regression behavior under-studied. This paper presents TDAD (Test-Driven Agentic Development), an open-source tool and benchmark methodology that combines abstract-syntax-tree (AST) based code–test graph construction with weighted impact analysis to surface the tests most likely affected by a proposed change. Evaluated on SWE-bench Verified with two local models (Qwen3-Coder 30B on 100 instances and Qwen3.5-35B-A3B on 25 instances), TDAD’s GraphRAG workflow reduced test-level regressions by 70% (6.08% $\to$ 1.82%) and improved resolution from 24% to 32% when deployed as an agent skill. A surprising finding is that test-driven development (TDD) prompting alone increased regressions (9.94%), revealing that smaller models benefit more from contextual information (which tests to verify) than from procedural instructions (how to do TDD). I additionally demonstrate an autonomous auto-improvement loop that raised resolution from 12% to 60% on a 10-instance subset with 0% regression. These findings suggest that for AI agent tool design, surfacing contextual information outperforms prescribing procedural workflows. All code, data, and logs are publicly available.
@@ 1. Introduction
Large-language-model (LLM) based coding agents have demonstrated impressive ability to resolve real-world software issues. On SWE-bench Verified (jimenez2024swebench, ), state-of-the-art agents resolve over 70% of GitHub issues from popular open-source repositories. Evaluation, however, centers almost exclusively on resolution rate: the fraction of issues for which the agent’s patch passes the issue-specific tests. A complementary question, how many previously-passing tests does the patch break?, receives far less attention.
Regressions matter in practice. A patch that fixes one bug but breaks three others would be rejected in any code-review process. Yet detecting regressions is fundamentally hard for agents without dependency awareness. An agent faces two options: (a) run all tests in the repository, which does not scale to large codebases where test suites take minutes or hours, or (b) run only the tests near the changed files, which misses indirect dependencies where code several modules away breaks due to a shared interface change. Diffs alone do not reveal these transitive relationships. In the baseline experiments, a vanilla agent caused 562 pass-to-pass (P2P) test failures across 100 instances, an average of 6.5 broken tests per generated patch. The severity varies widely: on astropy-13977, a single patch broke all 322 P2P tests, while on django-13089 TDD prompting alone turned 4 failures into 352, a complete regression of the test suite. These are not edge cases; they represent the default behavior of capable agents operating without structural awareness of code–test dependencies. This paper introduces that awareness.
The standard SWE-bench evaluation harness already collects the data needed to measure regressions: the PASS_TO_PASS test set records tests that passed before the gold patch and should still pass afterward. Yet this data is not surfaced in leaderboard rankings or agent evaluation reports. Recent evidence reinforces this concern: METR (metr2026swebench, ) found that roughly half of SWE-bench-passing patches would not be merged by real maintainers, and Ehsani et al. (ehsani2026agentfail, ) show that CI/CD failures are a leading cause of rejected agent-authored pull requests in practice. I argue that regression rate deserves first-class status alongside resolution rate.
This paper presents TDAD (Test-Driven Agentic Development), an open-source tool and benchmark methodology for measuring and reducing regressions in AI coding agents. TDAD builds an AST-derived code–test dependency graph, applies weighted impact analysis, and surfaces the results as a lightweight agent skill: a static test map and a 20-line instruction file. The design prioritizes zero-dependency integration: the agent needs only grep and pytest at runtime, with no graph database, MCP server, or API calls. The key insight is that agents do not need to be told how to do TDD; they need to be told which tests to check.
The key findings, evaluated on SWE-bench Verified with two local models:
@@
    70% regression reduction (100 instances, Qwen3-Coder 30B). TDAD reduced test-level regression rate from 6.08% to 1.82% (562 $\to$ 155 P2P failures).
@@
    TDD prompting paradox. TDD instructions without graph context increased regressions to 9.94%, worse than vanilla.
@@
    Resolution improvement as a skill (25 instances, Qwen3.5-35B-A3B + OpenCode agent). TDAD improved resolution from 24% to 32% and generation from 40% to 68%.
@@
    Autonomous self-improvement. An auto-improvement loop raised resolution from 12% to 60% (10-instance subset) with 0% regression.
@@
The remainder of this paper is organized as follows. Section 2 surveys related work on AI coding agents, regression testing, and graph-based code analysis. Section 3 describes TDAD’s architecture and algorithms. Section 4 details the experimental setup, and Section 5 presents results and analysis. Section 6 discusses implications and limitations, and Section 7 concludes with future directions.
@@ 2.2. Regression Testing
TDAD adapts these classical ideas to a novel setting: AI coding agents, where changes are generated programmatically and must be validated before submission. Unlike traditional RTS, which optimizes developer workflow, TDAD uses impact analysis to inform the agent about which tests are at risk, enabling self-correction before committing.
@@ 2.4. TDD and AI Agents
Test-driven development (TDD) (beck2003tdd, ) prescribes writing tests before implementation code, creating a tight feedback loop that catches regressions early. Cui (cui2025testsasprompt, ) introduces a TDD benchmark where test cases serve as both prompt and verification, finding that instruction following and in-context learning matter more than general coding proficiency for TDD success, and that performance degrades when instructions are long. This finding foreshadows the TDD prompting paradox observed in the present work (§5.2).
@@ 3.5. Agent Integration: TDAD as a Skill
TDAD integrates with agents through two static artifacts. test_map.txt maps source files to test files (one line per mapping, grep-able). SKILL.md is a 20-line definition: (1) fix the bug, (2) grep test_map.txt for related tests, (3) run them and fix failures. This brevity proved critical: through the auto-improvement loop (§5.5), I found that simplifying SKILL.md from 107 to 20 lines alone quadrupled resolution from 12% to 50%.
@@ 3.6. Backend Architecture
TDAD originally used Neo4j with Docker for graph storage. Through iterative development, I migrated to a NetworkX in-memory backend as the default, eliminating the Docker dependency entirely. Installation is now pip install tdad with zero external requirements. Graph persistence uses pickle (.tdad/graph.pkl). Neo4j remains available via TDAD_BACKEND=neo4j for large-scale deployments. Both backends expose the same GraphDB interface.
@@ 4.2. Models and Agents
Phase 1 (100 instances) uses Qwen3-Coder 30B (qwen2025qwen3coder, ) (Q4_K_M 4-bit quantization) served via llama.cpp on consumer hardware, with a 32K-token context window and temperature 0 for deterministic outputs. Each instance was given up to 15 minutes for patch generation. This phase evaluates regression reduction across three prompt configurations. I chose a smaller, locally-run model for three reasons: (1) to demonstrate that TDAD’s benefits are not contingent on frontier-scale models; (2) to enable fully reproducible experiments without API costs; and (3) to stress-test the approach where every context token matters.
@@ 5.1. Phase 1: Regression Reduction (100 Instances)
The three configurations produced different numbers of non-empty patches (86, 75, and 74 respectively), yielding different P2P test pools (9,245 vs. 8,040 vs. 8,536 total tests). The instance-level regression rates (30.2%, 33.3%, 33.3%) appear to contradict the test-level result; however, in absolute terms GraphRAG produced fewer regressing instances (25 of 74) than vanilla (26 of 86). The higher rate is a denominator artifact of GraphRAG’s lower generation rate. I report both metrics for transparency but emphasize test-level regression rate as primary because it captures regression severity: a patch breaking 322 tests is qualitatively different from one breaking 1.
@@ 5.4. Phase 2: TDAD as an Agent Skill (25 Instances)
After the Phase 1 findings and an auto-improvement loop (§5.5), I packaged TDAD as a reusable skill with a NetworkX backend (no Docker) and evaluated it with a different model and agent framework. Table 6 shows the results.
@@ 5.5. Auto-Improvement Loop
Inspired by Karpathy’s autoresearch framework, [^1] https://github.com/karpathy/autoresearch I built an outer loop that autonomously improves TDAD’s components through iterative agent-driven modification and evaluation.
@@ 6.2. Regression as a First-Class Metric
Current AI coding benchmarks create a perverse incentive: agents are rewarded for resolution regardless of collateral damage. In practice, a patch that resolves one issue but introduces three regressions has negative value. METR (metr2026swebench, ) demonstrates this gap concretely: when maintainers of scikit-learn, Sphinx, and pytest reviewed 296 SWE-bench-passing patches, roughly half would not have been merged, with regression and code quality among the top rejection reasons. I advocate for reporting regression rate alongside resolution rate as standard practice. The data is already collected by the SWE-bench Docker harness; it simply needs to be surfaced in evaluation reports and leaderboard rankings.
@@ 6.3. Limitations
  • Scale and statistical power. 100 instances (Phase 1) and 25 instances (Phase 2) without formal significance tests. The Phase 1 effect size is large (72% reduction, 562 vs. 155 failures), suggesting practical significance, but smaller effects could be noise at this sample size. Results should be confirmed on the full 500-instance set.
@@ 6.4. Threats to Validity
Configurations were run sequentially rather than interleaved, so time-dependent factors (e.g., thermal throttling on consumer hardware) could systematically favor earlier runs. This risk is mitigated by temperature 0 for deterministic decoding and Docker isolation for evaluation. The auto-improvement loop used a fixed 10-instance evaluation set; overfitting to those instances cannot be ruled out. However, two factors mitigate this concern: (1) the accepted changes were structural improvements to test-linking heuristics (directory proximity, import-based fallback) and prompt simplification, none of which encode instance-specific knowledge; and (2) the Phase 2 evaluation on a disjoint 25-instance set with a different model and agent confirmed that the optimized tool generalizes beyond the loop’s training set.
@@ 7. Conclusion
Future work includes extending TDAD to multiple languages via Tree-sitter, integrating dynamic coverage data to improve impact precision, and evaluating with frontier models (e.g., Claude Opus 4.6, GPT-5.4) to test whether the TDD prompting paradox persists at larger scale. I advocate for the community to adopt composite metrics that jointly capture resolution and regression, ensuring agents are evaluated on net contribution.
@@ 8. Data Availability
All artifacts are publicly available:
  • Install: pip install tdad (zero dependencies beyond NetworkX)
  • Experiment log: EXPERIMENTS.md (28 experiments, EXP-001 through EXP-028)
  • All patches, evaluation logs, and auto-loop history archived in the repository under MIT license.

### 19 Mar 2026 v2

@@ Abstract.
AI coding agents can resolve real-world software issues, yet they frequently introduce regressions—breaking tests that previously passed. Current benchmarks focus almost exclusively on resolution rate, leaving regression behavior under-studied.
This work develops a method to enhance agents with appropriate structural knowledge that leads to meaningful regression reduction while improving resolution rate, and presents TDAD (Test-Driven Agentic Development), an open-source tool that performs pre-change impact analysis for AI coding agents. TDAD builds a dependency map between source code and tests so that before committing a patch, the agent knows which tests to verify and can self-correct. The map is delivered as a lightweight agent skill—a static text file the agent queries at runtime. Evaluated on SWE-bench Verified with two open-weight models running on consumer hardware (Qwen3-Coder 30B, 100 instances; Qwen3.5-35B-A3B, 25 instances), TDAD reduced regressions by 70% (6.08% $\to$ 1.82%) compared to a vanilla baseline. In contrast, adding TDD procedural instructions without targeted test context increased regressions to 9.94%—worse than no intervention at all. When deployed as an agent skill with a different model and framework, TDAD improved issue-resolution rate from 24% to 32%, confirming that surfacing contextual information outperforms prescribing procedural workflows. All code, data, and logs are publicly available.
@@ 1. Introduction
Large-language-model (LLM) based coding agents have demonstrated impressive ability to resolve real-world software issues. On SWE-bench Verified (jimenez2024swebench, ), state-of-the-art agents resolve over 70% of GitHub issues from popular open-source repositories. Evaluation, however, centers almost exclusively on resolution rate: the fraction of issues for which the agent’s patch passes the issue-specific tests. A complementary question—how many previously-passing tests does the patch break?—receives far less attention, despite recent evidence that regressions and CI/CD failures are a leading cause of rejected agent-authored pull requests (metr2026swebench, ; ehsani2026agentfail, ).
Regressions are hard to prevent without dependency awareness. An agent can either run all tests (too slow for large codebases) or run only tests near the changed files (missing indirect dependencies). This problem parallels classical regression test selection (RTS) (elbaum2014techniques, ; legunsen2016extensive, ), but the agentic context introduces novel requirements: changes are generated programmatically, the agent has a limited context window, and test verification must happen before submission rather than in a CI pipeline afterward. In the baseline experiments, a vanilla agent caused 562 pass-to-pass (P2P) test failures across 100 instances—an average of 6.5 broken tests per generated patch.
The SWE-bench evaluation harness already collects the data needed to measure regressions: the PASS_TO_PASS test set records tests that passed before the gold patch and should still pass afterward. Yet this data is not surfaced in leaderboard rankings. Recent evidence reinforces this concern: METR (metr2026swebench, ) found that roughly half of SWE-bench-passing patches would not be merged by real maintainers, and Ehsani et al. (ehsani2026agentfail, ) show that CI/CD failures are a leading cause of rejected agent-authored pull requests. Regression rate deserves first-class status alongside resolution rate.
This paper presents TDAD (Test-Driven Agentic Development), an open-source tool that performs pre-change impact analysis for AI coding agents. TDAD builds a dependency map between source code and tests so that before committing a patch, the agent knows exactly which tests to verify and can self-correct (Figure 1). The key insight is that agents do not need to be told how to do TDD; they need to be told which tests to check. At runtime the agent needs only grep and pytest—no graph database, MCP server, or API calls.
The evaluation spans two phases on SWE-bench Verified—a controlled comparison on 100 instances with a single model, followed by a generalization study with a different model and agent framework—yielding four key findings:
@@
    70% regression reduction. In Phase 1 (100 instances, Qwen3-Coder 30B, single-agent setup), TDAD reduced test-level regression rate from 6.08% to 1.82% (562 $\to$ 155 P2P failures).
@@
    TDD Prompting Paradox. In the same Phase 1 setup, an ablation adding TDD procedural instructions (write tests first, then implement) without telling the agent which specific tests to check actually increased regressions to 9.94%—worse than vanilla. Procedural instructions without targeted context are counterproductive.
@@
    Resolution improvement as agent skill. In Phase 2 (25 instances, different model Qwen3.5-35B-A3B + OpenCode agent), TDAD improved resolution from 24% to 32% and generation from 40% to 68%, confirming generalization beyond the original setup.
@@
    Autonomous self-improvement. An auto-improvement loop (10-instance subset) iteratively refined TDAD’s own configuration, raising resolution from 12% to 60% with 0% regression.
ReadissuePlanfixImplementTDAD:at-risk testsVerifytestsSubmitpatchfix regressions
Figure 1. Agentic workflow with TDAD. After implementing changes, the agent queries TDAD’s test map to identify at-risk tests, verifies them, and self-corrects if regressions are detected. The dashed arrow represents the self-correction loop.
@@ 2.2. Regression Testing
TDAD adapts these classical RTS ideas to a novel setting: AI coding agents. The differences are threefold. Timing: traditional RTS optimizes CI time by selecting which tests to run after a change is committed; TDAD operates before the agent commits, enabling self-correction rather than post-hoc detection. Consumer: the output of traditional RTS feeds a test runner or CI pipeline; TDAD’s output is a static text file designed to be consumed by an LLM agent via grep, respecting context-window limits. Goal: traditional RTS minimizes execution time for a fixed test suite; TDAD minimizes regressions by surfacing which tests the agent should verify, not merely which tests to run.
@@ 2.4. TDD and AI Agents
Test-driven development (TDD) (beck2003tdd, ) prescribes writing tests before implementation code, creating a tight feedback loop that catches regressions early. TDAD borrows TDD’s core philosophy—verify before you commit—but replaces TDD’s procedure (write test $\to$ red $\to$ green $\to$ refactor) with contextual guidance: rather than instructing the agent to follow TDD steps, TDAD tells the agent which specific tests are at risk. This distinction directly motivates the TDD Prompting Paradox (§5.2), where procedural TDD instructions without contextual information actually increased regressions. Cui (cui2025testsasprompt, ) introduces a TDD benchmark where test cases serve as both prompt and verification, finding that instruction following and in-context learning matter more than general coding proficiency for TDD success, and that performance degrades when instructions are long—a result consistent with the present work’s finding that smaller models benefit more from concise, targeted context than from lengthy procedural instructions.
@@ 3.4. Impact Analysis
Justification of constants. The confidence weight $c_{w}=0.3$ allocates 70% of the score to strategy-specific weights and 30% to link-quality confidence, ensuring that the type of relationship (direct test, transitive call chain, coverage, import) dominates while still rewarding stronger evidence. The confidence values themselves—1.0, 0.56, 0.5, 0.45—form a monotonically decreasing scale reflecting trust in link quality: direct TESTS edges are ground-truth associations, transitive call chains decay with hop distance, file-level coverage is coarser, and import-level links are the weakest signal. These values were set heuristically from the graph structure and subsequently refined through the autonomous auto-improvement loop (§5.5), where the agent itself iterated on weight configurations to minimize regressions. The three weight profiles (conservative, balanced, aggressive) allow practitioners to adapt to different codebases without modifying the formula.
@@ 3.5. Agent Integration: TDAD as a Skill
TDAD integrates with agents through two static artifacts. test_map.txt maps source files to test files (one line per mapping, grep-able). SKILL.md is a 20-line definition: (1) fix the bug, (2) grep test_map.txt for related tests, (3) run them and fix failures. This brevity proved critical: through the auto-improvement loop (§5.5), we found that simplifying SKILL.md from 107 to 20 lines alone quadrupled resolution from 12% to 50%.
@@ 3.6. Backend Architecture
TDAD originally used Neo4j with Docker for graph storage. Through iterative development, we migrated to a NetworkX in-memory backend as the default, eliminating the Docker dependency entirely. Installation is now pip install tdad with zero external requirements. Graph persistence uses pickle (.tdad/graph.pkl). Neo4j remains available via TDAD_BACKEND=neo4j for large-scale deployments. Both backends expose the same GraphDB interface.
@@ 4.2. Models and Agents
Phase 1 (100 instances) uses Qwen3-Coder 30B (qwen2025qwen3coder, ) (Q4_K_M 4-bit quantization) served via llama.cpp on consumer hardware, with a 32K-token context window and temperature 0 for deterministic outputs. Each instance was given up to 15 minutes for patch generation. This phase evaluates regression reduction across three prompt configurations. We chose a smaller, locally-run model for three reasons: (1) to demonstrate that TDAD’s benefits are not contingent on frontier-scale models; (2) to enable fully reproducible experiments without API costs; and (3) to stress-test the approach where every context token matters.
@@ 5.1. Phase 1: Regression Reduction (100 Instances)
The three configurations produced different numbers of non-empty patches (86, 75, and 74 respectively), yielding different P2P test pools (9,245 vs. 8,040 vs. 8,536 total tests). The instance-level regression rates (30.2%, 33.3%, 33.3%) appear to contradict the test-level result; however, in absolute terms GraphRAG produced fewer regressing instances (25 of 74) than vanilla (26 of 86). The higher rate is a denominator artifact of GraphRAG’s lower generation rate. We report both metrics for transparency but emphasize test-level regression rate as primary because it captures regression severity: a patch breaking 322 tests is qualitatively different from one breaking 1.
@@ 5.4. Phase 2: TDAD as an Agent Skill (25 Instances)
After the Phase 1 findings and an auto-improvement loop (§5.5), we packaged TDAD as a reusable skill with a NetworkX backend (no Docker) and evaluated it with a different model and agent framework. Table 6 shows the results.
@@ 5.5. Auto-Improvement Loop
Inspired by Karpathy’s autoresearch framework, [^1] https://github.com/karpathy/autoresearch we built an outer loop that autonomously improves TDAD’s components through iterative agent-driven modification and evaluation.
@@ 6.2. Regression as a First-Class Metric
Current AI coding benchmarks create a perverse incentive: agents are rewarded for resolution regardless of collateral damage. In practice, a patch that resolves one issue but introduces three regressions has negative value. METR (metr2026swebench, ) demonstrates this gap concretely: when maintainers of scikit-learn, Sphinx, and pytest reviewed 296 SWE-bench-passing patches, roughly half would not have been merged, with regression and code quality among the top rejection reasons. We advocate for reporting regression rate alongside resolution rate as standard practice. The data is already collected by the SWE-bench Docker harness; it simply needs to be surfaced in evaluation reports and leaderboard rankings.
@@ 6.3. Limitations
  • Scale and statistical power. 100 instances (Phase 1) and 25 instances (Phase 2) without formal significance tests. The Phase 1 effect size is large (72% reduction, 562 vs. 155 failures), suggesting practical significance, but smaller effects could be noise at this sample size. The 100-instance scope reflects hardware and time constraints of this initial research: all experiments ran on consumer hardware with locally-hosted models, where each instance requires several minutes of inference; a full 500-instance run would take multiple days per configuration. Results should be confirmed on the full 500-instance set as resources permit.
@@ 6.4. Threats to Validity
Configurations were run sequentially rather than interleaved, so time-dependent factors could systematically favor earlier runs; this is mitigated by temperature 0 and Docker isolation. The auto-improvement loop’s fixed 10-instance evaluation set risks overfitting, but the accepted changes were structural (test-linking heuristics, prompt simplification) and generalized to a disjoint 25-instance Phase 2 set.
@@ 7. Conclusion
Future work includes extending TDAD to multiple languages via Tree-sitter, integrating dynamic coverage data to improve impact precision, and evaluating with frontier models (e.g., Claude Opus 4.6, GPT-5.4) to test whether the TDD prompting paradox persists at larger scale. We advocate for the community to adopt composite metrics that jointly capture resolution and regression, ensuring agents are evaluated on net contribution.
@@ 8. Data Availability
All code, patches, evaluation logs, and 28 experiment records are publicly available at https://github.com/pepealonso95/TDAD under MIT license. Install via pip install tdad (zero dependencies beyond NetworkX).
@@
Acknowledgements
Sergio Yovine was partially funded by grant ANII FMV_1_2023_1_175864, Uruguay.